# Assortment Optimization Under History-Dependent Effects

### v2

@@ 7 Conclusion
This paper studies a multi-period assortment planning problem in which customer utility depends on historical assortments. We first show that the problem is solvable in polynomial time if and only if customers’ utility increases with historical repeated offerings. The optimal policy is a sequential revenue-ordered policy. However, the problem becomes NP-hard if the historical effect is the opposite, even with a planning horizon of two periods. We develop an MIECP formulation to find a global optimal solution. Our numerical analyses indicate that our formulation efficiently achieves optimality and generates higher revenue than heuristic policies, especially under strong satiation effects. We further explore cyclic policies when the problem has a large time horizon. We theoretically prove that a cyclic policy with a cycle equal to the memory length plus one is asymptotically optimal. Additionally, we develop a neat, bound-free reformulation for non-overlapping cyclic policies when the memory length equals one.
Our research can be extended in several directions. The first direction would be to explore the joint assortment and pricing or inventory problem under history-dependent utility. The second direction would be to determine, for online settings under the history-dependent effect, how to provide personalized assortments when customers stochastically revisit the online store. Third, it would be interesting to examine how to incorporate history-dependent effects into different choice models and optimize assortment decisions.

### v3

@@ 7 Conclusion
This paper studies a multi-period assortment planning problem with history-dependent customer utility. We prove that the problem is NP-hard under negative history-dependent effects. To address this for general history-dependent effects, we develop an MIECP formulation to find a global optimal solution. We also prove the optimality of the sequential policy for the addiction case and the asymptotic optimality of cyclic policies while relating the cycle length to customer memory length.
Our research can be extended in several directions. First, it is important to develop efficient approximation algorithms, such as fully polynomial-time approximation schemes, to address the history-dependent assortment planning problem, especially in cases with large memory lengths and long planning horizons. Second, extending the model to online settings for personalized assortments or dynamic settings to adjust assortments to quickly respond to stochastic demand. Third, investigating joint assortment, pricing, or inventory problems under history-dependent effects. Finally, exploring history-dependent effects in other choice models for assortment optimization.